OpenStudy (anonymous):

I have another question A square piece of paper is folded in half. the perimeter of the resulting rectangle is 24cm. find the area of the original square paper. Please help

OpenStudy (anonymous):

24+6=30..Original perimeter 30cm

OpenStudy (anonymous):

ow I got this answer x+x+x/2+x/2 = 3x 3x=24 x=8 The area of originalpaper is 8*8 = 64

OpenStudy (anonymous):

 In the drawing you will find the smaller square with sizes x and the 'folded out' larger rectangle with sizes 2x by x. For both I have incorporated the calculations of the Area (A) and the Circumference (C). Only one of these is given: the circumference of the square (C1), which is the same as the perimeter ofcourse. You can solve C1 now because you know C1 = 24, which allows you to calculate x. Once found, you can calculate A2 by filling in x into the equation. Do you know how to do that ?

OpenStudy (anonymous):

ow